# Words as Trigger Points in Social Media Discussions: A Large-Scale Case Study about UK Politics on Reddit

### v3

@@ 5 Conclusion
This paper establishes trigger words as a concept in online debates. Since our work represents the first systematic computational study of this concept, our research design and case selection are geared to clearly identify such behaviour in the most likely cases. To do so, we performed a large-scale analysis about UK politics on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS, and Vaccine).
@@
The results show that trigger words cause more user engagement, and this activity is marked by higher levels of animosity. This confirms the first hypothesis (H1), which suggests more user interaction after the appearance of a trigger word. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity of the ensuing discussions, we examine four different features that can indicate such effects. Trigger words cause more controversiality, more negative sentiment, and more anger. We also investigate the causal effect of trigger words on hate speech. Where we can, we rely on a model capable of identifying domain-specific forms of hate —sexism in the case of Feminism and racism in the case of Rwanda—. Our findings similarly confirm that trigger words play a causal role in this aspect as well.
@@
From this work, we can certify that trigger points exist in online communication, and we have provided a framework to measure the extent of them in terms of user engagement and animosity. Future work will analyse this phenomenon more comprehensively. We aim to broaden the scope, study how prevalent trigger points are in different countries and languages, and seek to better understand further behavioural consequences from trigger points. It is our intention to evaluate the extent to which trigger points are responsible for hateful online communication. We also strive to automate the trigger point detection process to detect the respective words in online conversations when we do not know them ex-ante.
6 Limitations
@@
We refrain from using large language models (LLMs) for feature extraction mainly due to the vast size of our dataset ($>$100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, existing literature indicates that the zero-shot capabilities of LLMs do not consistently outperform smaller models for specific social media tasks like ours (Antypas et al. 2023).
@@
Our study does not attempt to map how pervasive trigger points are in online debates. The chosen subreddits may not fully represent the broader range of discussions these trigger words could provoke across the diverse Reddit community. Similarly, we acknowledge the limitations and potential biases of our work, as we only investigate a small set of potential trigger words in the context of the UK and limit our data to comments written in English in a limited selection of subreddits. The choice of trigger words is informed by a contextual analysis of the UK politics, and the Reddit threads have been selected due to the often polarised conversations they host around such topics. However, this is the first study that analyses trigger points’ concept validity in online conversations. Based on our empirical evidence, we can be confident that we have found relevant evidence that warrants further investigation.
7 Ethics Statement
@@
All data used in our research is sourced from publicly available information or collected using the official Reddit API. It is provided in an aggregated way and without reporting sensitive information from individual users.
@@
The automatic methods used can be abused by those with the power to repress people with opposing opinions. Hence, we offer a comprehensive study but do not share our data publicly to avoid potential misuse or commercial use. Researchers interested in replicating our experiments can contact us directly.
@@ Appendix A Model Selection
A.1 Emotion Categories
The twitter-roberta-base-emotion-multilabel model classifies each entry in one or more of the following classes: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust.
A.2 Hate Speech Categories.
The twitter-roberta-base-hate-multiclass model classifies each entry in one of the following classes: not_hate, sexism, racism, religion, other, sexual_orientation, disability
A.3 Validation Setting
Table 6 displays the distribution of comments in regards to emotion, sentiment and hate speech for the manually annotated dataset collected.
@@ A.3 Validation Setting
Table 6: Distribution of comments regarding emotion, sentiment, and hate speech categories on the manually annotated dataset.
A.4 Annotation Guidelines
Below we present the guidelines presented to annotators for the Sentiment Analysis, Emtion classification, and Hate Speech detection tasks.
Sentiment Analysis
We will show you Reddit comments that can be positive, negative, or neutral based on their sentiment.
Sentiment Categories
  • Positive comments should be labelled: Positive
  • Neutral comments should be labelled: Neutral
  • Negative comments should be labelled: Negative
Examples
  • “‘Just finished my morning workout and feeling fantastic! #HealthyLiving #FitnessGoals“‘ is a positive sentence.
  • “‘Stuck in traffic again! Late for work and already stressed out. #TrafficJam“‘ #MondayBlues is a negative sentence.
  • Factual sentences such as “‘The party was in front of the house“‘ are neutral sentences.
Notes: We have also added a comment column where you can share your thoughts, comments, or doubts. If you are unsure whether you should annotate the tweet as positive, negative or neutral, you may choose the label that suits it best and a comment about that in the comment column.
The annotation of an ironic comments depends on the context. While we acknowledge the lack of context, if you think the tweet is mocking someone (or a group of people), you should label it negative.
Emotion Classification
You will be presented with Reddit comments and asked to select the option that best describes the text. You may choose as many emotions as you believe are present — from 0 to 11 — based on the content and tone of the comment.
Emotion Categories:
  • Anger – includes annoyance, rage
  • Anticipation – includes interest, vigilance
  • Disgust – includes disinterest, dislike, loathing
  • Fear – includes apprehension, anxiety, terror
  • Joy – includes serenity, ecstasy
  • Love – includes affection
  • Optimism – includes hopefulness, confidence
  • Pessimism – includes cynicism, no confidence
  • Sadness – includes pensiveness, grief
  • Surprise – includes distraction, amazement
  • Trust – includes acceptance, liking, admiration
Notes: If no emotion is clearly present, you may leave all options unchecked. Choose all emotions that are clearly expressed or implied in the comment. If you’re unsure, use the comment column to share your reasoning or highlight ambiguity. Comments may express multiple emotions simultaneously, for example, both sadness and anger.
Hate Speech Detection
You will be presented with Reddit comments and asked to identify whether they contain hate speech, and if so, which target group is being attacked or discriminated against. Please choose only one of the following categories that best represents the nature of the content:
Hate Speech Categories:
  • Sexism: Content that targets or demeans people based on their gender (e.g., women, men, non-binary individuals) or reinforces harmful gender stereotypes. Example: “Women should stay in the kitchen where they belong.”
  • Racism: Content that expresses hostility, prejudice, or dehumanization based on race, ethnicity, or national origin. Example: “All [ethnic group] are criminals and should be deported.”
  • Disability: Content that mocks, belittles, or degrades people with physical, mental, or developmental disabilities. Example: “You’re so stupid, you must be on the spectrum.”
  • Sexual Orientation: Hate directed toward someone because of their actual or perceived sexual orientation (e.g., gay, lesbian, bisexual, etc.). Example: “Being gay is a disease that should be cured.”
  • Religion: Disparagement, dehumanization, or threats against individuals or groups based on religious beliefs or identities. Example: “All Muslims are terrorists.”
  • Other: Use this label for hate speech that targets a group not covered above, such as socioeconomic status, age, appearance, political affiliation, etc. Example: “Old people are useless and should just die already.”
  • Not Hate: Use this label if the comments does not contain hate speech. Comments that are neutral or non-discriminatory.
Notes: If you are unsure, select the label you think fits best and add a comment explaining your reasoning. A single post may mention multiple groups in such a case, select the one that is most explicit or harmful, and explain in the comment column.
@@ Appendix B Extended Results
Table 7 presents the total number of comments and threads extracted from Reddit for each keyword, along with the corresponding time periods during which the data was collected.
@@
Figure 8 displays the number of comments posted before and after the introduction of each trigger point on the data collected.
Figure 9 displays the average controversiality per thread before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
Figure 10 displays the distribution of comments posted before and after the introduction of the Rwanda trigger point in the Temporal and Contextual control settings.
Figure 11 displays the distribution of negative comments, based on their sentiment, posted before and after the introduction of the Brexit, Feminism, Rwanda, and NHS trigger points in the Space control setting.
Figure 12 displays the distribution of comments conveying anger and those that do not before and after the introduction of the Vaccine trigger point in the Space control setting.
Threads Comments Year
Rwanda 0.05 1.87 2020, 2022
Brexit 0.37 11.57 2020 - 2022
NHS 0.34 6.20 2020 - 2022
Vaccine 1.42 28.11 2019, 2022
Feminism 0.94 17.79 2020 - 2022
EU 0.2 12.4 2020 - 2022
Red Cross 0.02 0.2 2020 - 2022
SubSaharan 0.08 1.5 2020, 2022
Surgery 0.5 3.9 2019, 2022
Woman 2.1 19.3 2020 - 2022
Table 7: Numbers, in millions, of threads, comments, and time period for each trigger word.
@@
Table 8: Number of comments present for each trigger in each experiment for both the treatment and control subsets.
(a) Brexit
(b) Feminism
(c) NHS
(d) Rwanda
(e) Vaccine
Figure 8: Absolute count of comments before and after each trigger point.
(a) Temporal Setting
(b) Contextual Setting
Figure 9: Average controversiality per thread in the Temporal and Contextual settings and their respective treatment groups, for the trigger point Rwanda.
@@
Figure 10: Rwanda Treatment and Control distributions of comments in the Contextual and Temporal settings.
(a) Brexit
(b) Feminism
(c) Rwanda
(d) NHS
Figure 11: Distribution of negatively charged comments in the Treatment and Control (Space) groups.
Figure 12: Distribution of anger and non-anger comments for the treatment and control (space) groups of Vaccine trigger word.
Appendix C Extended Test of Parallel Trends Assumption
Figures 13 a & b display the results of the Parallel Trends Assumption testing for all experiments as described in H1 and H2.
(a) Experiments from H1.
(b) Experiments from H2.
Figure 13: Testing the Parallel Trends Assumption with event studies. Individual studies in grey, their time-wise mean in purple.
Appendix D Computational Resources
An NVIDIA GeForce RTX 4090 GPU was utilised for the experiments conducted. 170 hours spent for the inference process of sentiment analysis, emotion detection, and hate speech detection models.
Appendix E Selected Reddit Comments
Below we present a Reddit discussion and how it evolves when the trigger ”feminism” appears. The comment containing the trigger is bolded.

### v4

@@ 6. Conclusion
Our work represents the first systematic computational study of trigger words in online debates. The research design and case selection are geared toward identifying such behaviour in the most likely cases, namely a large-scale analysis of UK politics on Reddit using five trigger words (Rwanda, Feminism, Brexit, NHS, and Vaccine).
@@
The results show that trigger words lead to more user engagement and higher levels of animosity. This confirms the first hypothesis (H1), which predicts increased user interaction upon encountering a trigger word. Regarding our second hypothesis (H2), which concerns the impact of trigger words on the animosity in ensuing discussions, we examine four features that may indicate such effects. Trigger words cause more controversiality, more negative sentiment, and more anger. Our findings also confirm that trigger words cause more hate speech.
@@
From this work, we can certify that trigger words exist in online communication. Future work will analyse this phenomenon more comprehensively. We aim to broaden the scope, study the prevalence of trigger words across countries and languages, and better understand their behavioural consequences. We also strive to automate trigger-word detection to identify relevant words in online conversations when we are not familiar with them a priori.
7. Limitations
@@
We refrain from using large language models (LLMs) for feature extraction primarily because of the size of our dataset (>>100 million entries). Using LLMs would have significantly increased the computational and financial cost and the ecological impact of our experiments. Moreover, the existing literature indicates that LLMs’ zero-shot capabilities do not consistently outperform smaller models on specific social media tasks such as ours (Antypas et al., 2023).
@@
Our study does not aim to map the prevalence of trigger words across online debates. The selected subreddits may not fully represent the broader range of discussions these trigger words could elicit across the Reddit community. Similarly, we acknowledge the limitations and potential biases of our work, as we investigate only a small set of trigger words in the UK and limit our data to English-language comments from a limited selection of subreddits. However, as the first study to analyse the validity of trigger words in online discussions, we are confident that our findings warrant further investigation.
8. Ethics Statement
@@
All data used in our research is sourced from publicly available information or collected using the official Reddit API. We present aggregated data without reporting sensitive information from individual users.
@@
The automatic methods used can be abused by those with the power to repress people with opposing opinions. Hence, we provide a comprehensive study but do not share our data publicly to prevent potential misuse or commercial exploitation. Researchers interested in replicating our experiments are encouraged to contact us directly.
@@ Appendix A Model Selection
A.1. Validation of Annotation Model
We use computationally efficient models that can be deployed locally to extract signals relevant to our hypotheses, and then validate the annotation procedure. We sampled 100 Reddit comments from our dataset, 20 for each trigger word, and then hand-annotated sentiment, emotion, and the presence of hate speech with three experienced NLP and CSS researchers. Each comment was assigned a label by majority vote, and disagreements were resolved through discussion (see Table 5). Given the difficulty of the annotation task, the annotations above show above-average agreement scores for sentiment and hate speech, with Fleiss’ $\kappa$ scores of 0.52 and 0.56, respectively. For the multi-label emotion annotation task, where annotators had to choose from a set of 12 labels, the Krippendorff’s alpha score (Krippendorff, 2011) reached 0.24, which is in line with other similar multi-label tasks with a high number of labels (Mohammad et al., 2018; Antypas et al., 2022; Ousidhoum et al., 2019).
@@ A.1. Validation of Annotation Model
Table 5. Distribution of comments regarding emotion, sentiment, and hate speech categories on the manually annotated dataset.
We fine-tune the base version of BERT-cased, XLNet, and RoBERTa for sentiment analysis, emotion recognition, and hate speech detection on the same datasets as the models we used for our experiments (Rosenthal et al., 2017; Mohammad et al., 2018; Antypas and Camacho-Collados, 2023).
We further compare our results against five publicly available pre-trained models on Hugging Face. Specifically, two sentiment classifiers trained on Reddit data777reddit-dBERT: https://huggingface.co/akshataupadhye/finetuning-sentiment-model-reddit-data, reddit-XLNET: https://huggingface.co/minh21/XLNet-Reddit-Sentiment-Analysis , hateBERT (Caselli et al., 2021) and one emotion classifier [^8] BERT-emotion: https://huggingface.co/ayoubkirouane/BERT-Emotions-Classifier. All models are evaluated in the aforementioned manually labelled dataset.
Table 6 presents the average micro F1-scores achieved by the models across each task. [^9] Micro F1-score provide a more realistic representation of the results, given the limited sample and heavily imbalanced data. The Twitter-RoBERTa models achieved the highest scores on all tasks. These scores are matched by XLNet-base for sentiment analysis and XLNet for emotion detection, and the fine-tuned models RoBERTa and XLNet for hate speech detection. Overall, the tested models performed comparably well, except for two off-the-shelf sentiment classifiers that achieved scores below 50%.
Sentiment Emotion Hate
t-RoBERTa 0.76 0.50 0.88
RoBERTa 0.75 0.47 0.88
XLNet 0.76 0.45 0.88
BERT 0.74 0.40 0.85
reddit-dBERT 0.45 - -
reddit-XLNet 0.31 - -
BERT-emotion - 0.45 -
hateBERT - - 0.80
Random 0.30 0.20 0.07
Table 6. Average micro-F1 scores of the models tested for the sentiment analysis, emotion, and hate speech detection tasks. Random indicates a naive baseline that randomly selects labels.
A.2. Annotated Categories
  • Emotion Categories: anger, anticipation, disgust, fear, joy, love, optimism, pessimism, sadness, surprise, trust
  • Hate Speech Categories: not_hate, sexism, racism, religion, other, sexual_orientation, disability
@@ Appendix B Extended Results
Figure 4 displays the distribution of comments posted before and after the introduction of the Rwanda trigger word in the Temporal and Contextual control settings.
@@
Table 7. Number of comments present for each trigger in each experiment for both the treatment and control subsets.
@@
Figure 4. Rwanda Treatment and Control distributions of comments in the Contextual and Temporal settings.
Appendix C Robustness of Results
Testing the Parallel Trends Assumption.
The DiD estimator relies on the assumption of parallel trends. While there are no formal guarantees that the assumption holds, there are grounds to argue in its favour. The first argument is theoretical. In neither of our experimental designs did we have reason to believe that the frequency of expressing animosity would not follow the same trend without the trigger word. Whether comparing treatment and control groups across different Reddit subreddits (spaces), different words (contextual), or points in time (temporal), without the trigger word, the relative difference between treatment and control groups would remain the same.
To further empirically support our argument, we conduct event studies. For each experiment, we model the outcome variable over time, enabling us to assess the dynamics of treatment impact before and after the intervention. Plotting the treatment effects at different time points before the treatment, the event study helps visualise whether the pre-treatment trends in the treatment and control groups are indeed parallel (Roth et al., 2023).
For each of our 60 overall experiments, we divide the 30 minutes before and after the first mention of a trigger word into six 10-minute intervals, rather than just two. This allows us to track pre-treatment trends at higher granularity. We calculate each discussion thread’s share of (hostile) messages in the six intervals. Averaging across all data points per period, we subtract the control group from the treatment group and compute the difference for each observation relative to the first observation period.
Figure 6 charts the time-wise means over all 60 event studies. If the parallel trends assumption holds, the lines before the treatment at time =0=0 should all remain on the horizontal line where the treatment effect =0=0. Indeed, this is essentially what we find. The mean treatment effect is close to 0 before the trigger word is first mentioned. After the intervention, it is clearly above the horizontal 0-line.
Figure 5. Placebo treatments every 10 minutes, ranging from 30 minutes before to 30 minutes after the actual treatment point.
Placebo Treatment.
We also implement a placebo treatment analysis. The core logic of this approach is to check whether the estimated treatment effects could be driven by other trends earlier or later than the actual intervention. Shifting the assumed treatment time in 10-minute intervals, ranging from 30 minutes before to 30 minutes after the actual intervention, we re-estimate all DiD models at each placebo point.
Figure 5 visualises these placebo tests by plotting the precision-weighted average treatment effects and their confidence intervals across all experiments measuring animosity (see Table 4). At the actual treatment point $t=0$, which compares message content in the 30 minutes before and after the first appearance of the trigger word ([30,0[[-30,0[ vs. $[0,30]$), we observe a clear and substantial effect (1.99). We then shift the treatment window rightward in 10-minute steps—first comparing [20,10[[-20,10[ to $[10,40]$, and so on—until the actual intervention fully exits the analysis window at $t=30$. This generates a steady decline in estimated effects, which converge to zero, as expected if no real treatment occurred at these placebo times. Similarly, shifting the window to earlier times (e.g., comparing [60,30[[-60,-30[ to [30,0][-30,0]) reveals a decreasing trend in effect size, though some residual difference remains. In conclusion, given the clear peak in the effect at $t=0$, we are confident in the robustness of our approach to defining the actual treatment moment.
Figure 6. Testing the Parallel Trends Assumption with event studies.
Appendix D Selected Reddit Comments
Below, we present a Reddit discussion and how it evolves when the trigger “feminism” is used. The comment containing the trigger is bolded.